OpenStudy (anonymous):

How should I solve this equation? 1/2(-6) - 2 (1/2 is a fraction >.<) Show steps please!

OpenStudy (anonymous):

\[\frac{ 1 }{ 2}-6-2\] or \[\frac{ 1 }{ 2}\times(-6)-2\]

OpenStudy (anonymous):

@Summer, it's the latter. Parentheses take precedence so the problem is solved linearly using (-6) as a complete term: \[\frac{ 1 }{ 2 }* (-6) = -3\] And then -3 -2 = -5